OpenStudy (anonymous):

Henry is 20 years older than George. In 5 years, Henry will be twice as old as George is then. Find their present ages.

OpenStudy (amistre64):

h = g+20 h+5 = 2g g+20+5 = 2g g+25 = 2g 25/2 = g ..... really? what i miss?

OpenStudy (amistre64):

h = g+20 h+5 = 2(g+5) maybe g+25 = 2g +10 15 = 2g ..... murmur

OpenStudy (amistre64):

gotta go; good luck ;)

OpenStudy (anonymous):

Henry is 45 and George is 25 The question states that: George is 20 younger than Henry (G=H-20) In five years, Henry will be 2 times as old as Gorge at present day (H+5=2G) Plug in G=H-20 into H+5=2G and solve for H then plug H (45) into the original formula to find that G is 25

OpenStudy (anonymous):

Thank you!